# Balancing Accuracy and Efficiency for Large-Scale SLAM: A Minimal Subset Approach for Scalable Loop Closures

### v2

@@ Abstract
Typical LiDAR SLAM architectures feature a front-end for odometry estimation and a back-end for refining and optimizing the trajectory and map, commonly through loop closures. However, loop closure detection in large-scale missions presents significant computational challenges due to the need to identify, verify, and process numerous candidate pairs for pose graph optimization. Keyframe sampling bridges the front-end and back-end by selecting frames for storing and processing during global optimization. This article proposes an online keyframe sampling approach that constructs the pose graph using the most impactful keyframes for loop closure. We introduce the Minimal Subset Approach (MSA), which optimizes two key objectives: redundancy minimization and information preservation, implemented within a sliding window framework. By operating in the feature space rather than 3-D space, MSA efficiently reduces redundant keyframes while retaining essential information. In sum, evaluations on diverse public datasets show that the proposed approach outperforms naive methods in reducing false positive rates in place recognition, while delivering superior ATE and RPE in metric localization, without the need for manual parameter tuning. Additionally, MSA demonstrates efficiency and scalability by reducing memory usage and computational overhead during loop closure detection and pose graph optimization.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) in large-scale missions presents significant challenges, particularly in maintaining accuracy and efficiency as the environment grows over time. A key component of back-end SLAM is Pose Graph Optimization (PGO), which ensures consistent mapping by optimizing the global trajectory. However, this process relies heavily on loop closure detection, to reduce accumulated drift, which is computationally expensive, especially in large-scale environments, due to the vast number of candidate pairs that need to be identified, verified, and processed to establish node connections in the pose graph. As the mission progresses, the number of loop closure candidates grows rapidly, leading to significant delays in processing times, with some datasets generating over 150,000 candidates, requiring four and a half hours to process, exceeding that of the mission time [1]. This complexity is further compounded by the computational intensity of data association and sequence matching during map optimization, where the processing time can range from seconds to several minutes, depending on the size of the data [2, 3]. Due to these computational challenges, many SLAM systems in large-scale environments have opted to either avoid or minimize loop closure detection to reduce the burden on limited computational resources. Some approaches, particularly in expansive environments, forego loop closures entirely, as the growing number of candidates becomes increasingly difficult to handle in real-time applications, as highlighted recently in [4]. While loop closure detection is critical for correcting drift and ensuring map consistency, the computational complexity can overwhelm systems if not managed effectively, often leading to false positives that degrade the overall quality of the map [5].
@@
Figure 1: An example on KITTI 06. Comparison of two trajectories from KISS-ICP, after Pose Graph Optimization (PGO). In the top figure, keyframes are sampled using an entropy-based approach, while in the bottom figure, keyframes are sampled with the proposed Minimal Subset Approach (MSA) [6] which achieves a lower Relative Pose Error (RPE).
I-A Related Work
Efficient implementations of pose graph optimization have significantly enhanced its scalability by leveraging factor graphs and incremental optimization techniques. In more detail, GTSAM [7] models probabilistic relationships using factor graphs and solves large-scale problems efficiently with solvers like iSAM2 [8], which incrementally updates the SLAM solution in real-time through a Bayes tree representation. Similarly, g2o [9] provides a versatile graph optimization framework, leveraging sparse pose graph connectivity for efficiency while supporting various parameterizations and loss functions. Furthermore, sparsification methods [10, 11, 12] and hierarchical approaches [13, 14] have further improved efficiency by reducing the number of active nodes in the optimization process. However, despite these advances, loop closure detection remains a major bottleneck due to the large number of candidate pairs that must be checked and processed to add meaningful constraints to the pose graph.
Keyframe sampling helps reduce the computational demands of loop closure detection by limiting the number of frames stored and processed in SLAM pipelines. In LiDAR-(Inertial) Odometry, systems such as LIO-SAM [15] and DiSCo-SLAM [16] or back-end global optimization systems like LAMP [17], use fixed Euclidean distance intervals (1–2 meters) for keyframe generation to maintain local maps and optimize global poses. Although effective, these methods lack adaptability in dynamic environments with varying densities of loop closure candidates. Recent advancements have introduced adaptive keyframe sampling techniques, adjusting intervals based on environmental spaciousness [18, 19], although manual threshold adjustment is still required. Entropy-based methods, like the one proposed by Zeng et al. [20], use information theory to select keyframes, but face similar adaptability issues. Other approaches, such as the one proposed by Ou et al. [21], focus on displacement vector similarity to balance computational cost and map completeness while managing keyframes in a sliding window.
@@ I-A Related Work
In contrast to the above that focus solely on back-end optimization, reducing the number of nodes in pose graph optimization, or generating keyframes for submap creation, our work centers on sampling keyframes online to reduce memory allocation and loop closure query times while preserving retrieval performance during place recognition and maintaining localization accuracy after pose graph optimization. In this work, we evaluate the impact of our previously introduced optimized keyframe sampling method [6] on critical back-end tasks, including loop closure detection and pose graph optimization, with a glimpse of the results illustrated in Fig. 1. The proposed Minimal Subset Approach (MSA) employs two key criteria, redundancy minimization and information preservation, within a sliding window optimization framework to streamline keyframe selection. By reducing redundant information while preserving essential data, MSA enhances scalability and computational efficiency for both place recognition and pose graph optimization. Unlike traditional methods that rely on spatial heuristics and 3-D space, MSA utilizes the feature space, making it inherently environment-agnostic. Its reliance on the descriptor extraction framework used for loop closure detection eliminates the need for manual parameter tuning, ensuring consistent performance across diverse scenarios.
I-B Contributions
Based on the aforementioned, the main contributions of this article can be summarized as follows: (a) To the best of our knowledge, this is the first paper to present the pose graph optimization (PGO) problem with an emphasis on the interplay between place recognition, keyframe sampling and loop closure detection. We highlight the importance of addressing these interconnected problems jointly to achieve a more comprehensive solution and provide further insights on the inherent challenges. (b) This paper significantly extends our previous work [6] by demonstrating the impact of keyframe sampling not only in place recognition but on pose graph optimization and metric localization as well. This provides valuable insights into the interconnection between the front-end and back-end SLAM, showcasing how sampling intervals influence the performance of the latter. (c) Through our results we demonstrate that MSA enhances scalability and efficiency by reducing computational demands. It consistently achieves lower false positive rates in place recognition and superior ATE and RPE in metric localization across diverse environments, without manual parameter tuning. By minimizing loop closure query times, memory usage, and the overall time required for pose graph optimization, MSA is well-suited for large-scale missions.
II Problem Formulation
II-A Pose Graph Optimization
@@
The objective of PGO is to find the most likely configuration of a robot’s trajectory by minimizing the error in a graph-based representation of the poses. The pose graph is represented as a set of nodes and edges, where each node represents a robot pose and each edge encodes a spatial constraint between two poses [22]. Let $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ be the set of $N$ robot poses, where, $\mathbf{x}\in SE(3)$, $\mathbf{z}_{ij}=h(\mathbf{x}_{i},\mathbf{x}_{j})$ be the observed spatial constraint between a pair of poses $\langle i,j\rangle$ and $\hat{\mathbf{z}}_{ij}(\mathbf{x}_{i},\mathbf{x}_{j})$ be the predicted value of that constraint. The error $\mathbf{e}_{ij}\in\mathbb{R}^{6}$ quantifies the difference between the observed relative transformation and the predicted transformation between poses $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ and is denoted as:
$\mathbf{e}_{ij}=\mathbf{e}(\mathbf{x}_{i},\mathbf{x}_{j})=\mathbf{z}_{ij}-\hat{\mathbf{z}}_{ij}(\mathbf{x}_{i},\mathbf{x}_{j})$ (1)
@@
We can then define the set of edges =$\mathcal{E}=\big{\{}\langle i,j\rangle\>|\>\exists\,\mathbf{z}_{ij}\big{\}}$, that contains all the pairs in which a relative constraint exists. The goal of pose graph optimization is to find the optimal configuration of poses $\mathbf{X}^{*}$ that minimizes the overall error function $E(\mathbf{X})$, denoted as:
@@
II-B Odometry and Loop Closure Edges
Pose graph optimization is widely used as a global optimization method to mitigate accumulated pose drift during large-scale and long-term missions. This would not be possible without loop closures. Relying solely on pose-to-pose odometry constraints is insufficient to maintain global consistency, hence, loop closures are essential. Given this, it is reasonable to assume that the set of edges $\mathcal{E}$ in the pose graph can be partitioned into two distinct subsets. The first subset, $\mathcal{E}_{o}\subseteq\mathcal{E}$, consists of odometry edges, which represent consecutive movements of the robot based on local motion estimates, while the second subset, $\mathcal{E}_{l}\subseteq\mathcal{E}$, consists of loop closure edges, which correspond to observations of previously visited locations. Thus, the overall error function can be decomposed into two components:
@@ II-B Odometry and Loop Closure Edges
II-C Loop Closure Detection
Detecting loop closures is challenging, particularly in large-scale missions, where accumulated drift can cause traditional radius-based searches to fail by missing nearby candidates. A common solution is place recognition, which compares current sensory data (e.g., images or LiDAR scans) against a database of past observations to identify matches. To formalize this, we define a keyframe set $\mathbf{K}=\{(\mathbf{x}_{i},\mathbf{d}_{i},\mathbf{z}_{i})\,|\,i=1,\ldots,N\}$, where $\mathbf{x}_{i}\in SE(3)$ is the pose of keyframe $i$, $\mathbf{d}_{i}\in\mathbb{R}^{M}$ represents its descriptor which encodes distinct features [23, 24], and $\mathbf{z}_{i}$ is the set of associated measurements. The goal is to find loop closure candidates $\mathcal{L}$ by comparing descriptors:
=$\mathcal{L}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}\times\mathbf{K},\;\mathbf{k}_{i}\neq\mathbf{k}_{j}\big{\}},$ (4)
@@ II-C Loop Closure Detection
II-D Reducing Search Space in Place Recognition
The detection and verification of loop closure edges $\mathcal{E}_{l}\triangleq\mathcal{L}$ is computationally demanding due to the large number of potential keyframe pairs, which scales quadratically as $O(n^{2})$, where $n=|\mathbf{K}|$ is the number of keyframes. While using descriptors and methods like k-nearest neighbor (k-NN) search [25] can accelerate the process, additional steps are still required to remove outliers, mitigate false positives, or prioritize certain nodes [1], while extracting the relative pose transformation between loop closure nodes also involves computationally intensive techniques such as General Iterative Closest Point (GICP) [26]. This pipeline does not scale well for extensive missions, often resulting in processing times of 3-4 hours or more [4, 17], and the challenge becomes even greater in multi-robot systems [3].
@@ II-D Reducing Search Space in Place Recognition
$\mathbf{K}_{\text{S}}^{*}=\operatorname*{arg\,min}_{\mathbf{K}_{\text{S}}\subseteq\mathbf{K}}\left(|\mathbf{K}_{\text{S}}|\right),\;\text{subject to}\;\mathcal{L}^{*}=\mathcal{L},\text{ where }$ (5)
=$\mathcal{L}^{*}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}^{*}_{\text{S}}\times\mathbf{K}^{*}_{\text{S}}\big{\}},$ (6)
@@ II-E Underlying Challenges
Addressing the combined problem of pose graph optimization and loop closure detection through place recognition presents several underlying challenges that affect both the efficiency and accuracy of the solution.
II-E1 Impact of Sampling
While the sampling strategy $S$ aims to preserve key loop closure candidates, in practice it must retain a subset $\bar{\mathcal{L}}\subseteq\mathcal{L}$ that still ensures an effective solution for pose graph optimization. Redundant loop closure edges can lead to issues, such as overfitting, increased computational burden, and numerical instability in the optimization process. Moreover, because the sampling process reduces the set of poses, the odometry edge set $\mathcal{E}_{o}^{*}\subset\mathcal{E}_{o}$ is also reduced, which can affect the overall accuracy of the optimization  [27], with an example illustrated on Fig. 2.
II-E2 Combined Problem
Combining loop closure detection through place recognition with pose graph optimization presents challenges due to their differing characteristics. Loop closure detection relies on matching high-dimensional feature descriptors, which is computationally expensive and sensitive to environmental factors. Conversely, pose graph optimization relies on geometric consistency and is affected by noise and outliers in loop closures. Balancing these require careful consideration of both feature reliability and constraint robustness to prevent false positives and reduced efficiency from degrading optimization performance.
II-E3 Dynamic Keyframe Selection

### v3

# A Minimal Subset Approach for Informed Keyframe Sampling in Large-Scale SLAM

@@ Abstract
Typical LiDAR SLAM architectures feature a front-end for odometry estimation and a back-end for refining and optimizing the trajectory and map, commonly through loop closures. However, loop closure detection in large-scale missions presents significant computational challenges due to the need to identify, verify, and process numerous candidate pairs for pose graph optimization. Keyframe sampling bridges the front-end and back-end by selecting frames for storing and processing during global optimization. This article proposes an online keyframe sampling approach that constructs the pose graph using the most impactful keyframes for loop closure. We introduce the Minimal Subset Approach (MSA), which optimizes two key objectives: redundancy minimization and information preservation, implemented within a sliding window framework. By operating in the feature space rather than 3-D space, MSA efficiently reduces redundant keyframes while retaining essential information. Evaluations on diverse public datasets show that the proposed approach outperforms naive methods in reducing false positive rates in place recognition, while delivering superior ATE and RPE in metric localization, without the need for manual parameter tuning. Additionally, MSA demonstrates efficiency and scalability by reducing memory usage and computational overhead during loop closure detection and pose graph optimization.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) in large-scale missions presents significant challenges, particularly in maintaining accuracy and efficiency as the environment grows over time. A key component of back-end SLAM is Pose Graph Optimization (PGO), which ensures consistent mapping by optimizing the global trajectory.
@@
Figure 1: An example on KITTI 06. Comparison of two trajectories from KISS-ICP, after Pose Graph Optimization (PGO). In the top figure, keyframes are sampled using an entropy-based approach, while in the bottom figure, keyframes are sampled with the proposed Minimal Subset Approach (MSA) which achieves a lower Relative Pose Error (RPE).
However, this process relies heavily on Loop Closure Detection (LCD) to reduce accumulated drift. In large-scale environments it becomes computationally expensive, as vast numbers of candidate pairs must be identified, verified, and processed to establish pose graph connections. As the mission progresses, the number of loop closure candidates grows rapidly, leading to significant delays in processing times. Some datasets generate over 150,000 candidates, requiring four and a half hours to process, exceeding that of the mission time [Denniston2022LoopCP]. This complexity is further compounded by the computational intensity of data association and sequence matching during map optimization, where the processing time can range from seconds to several minutes, depending on the size of the data [stathoulopoulos2024frame]. Due to these computational challenges, many SLAM systems in large-scale environments have opted to either avoid or minimize loop closure detection to reduce the burden on limited computational resources. Some approaches, particularly in expansive environments, forego loop closures entirely, as the growing number of candidates becomes increasingly difficult to handle in real-time applications, as highlighted recently by Ebadi et al. [Kamak2024darpa_slam]. While loop closure detection is critical for correcting drift and ensuring map consistency, the computational complexity can overwhelm systems if not managed effectively, often leading to false positives that degrade the overall quality of the map [cadena2016slam].
@@
In order to address these challenges, our work centers on sampling keyframes online to reduce memory allocation and loop closure query times while preserving retrieval performance during Place Recognition (PR) and maintaining localization accuracy after pose graph optimization. We evaluate the impact of our optimized keyframe sampling method on critical back-end tasks, including loop closure detection and pose graph optimization, with a glimpse of the results showing in Fig. 1. The proposed Minimal Subset Approach (MSA) employs two key criteria, redundancy minimization and information preservation, within a sliding window optimization framework to streamline keyframe selection. By reducing redundant information while preserving essential data, MSA enhances scalability and computational efficiency for both place recognition and pose graph optimization. Unlike traditional methods that rely on spatial heuristics in 3-D space, MSA operates in descriptor feature space. By leveraging the same extraction framework used for loop closure detection, it eliminates manual parameter tuning and achieves consistent performance across diverse scenarios.
Based on the aforementioned, the main contributions of this article can be summarized as follows: (a) To the best of our knowledge, this is the first paper to present a keyframe sampling methodology that jointly evaluates the impact of sampling on place recognition performance and on the localization accuracy of the resulting pose graph optimization. We emphasize addressing these components together and discuss the theoretical and practical challenges that arise from this coupled view. (b) We propose the Minimal Subset Approach (MSA), a keyframe sampling strategy that balances information preservation and redundancy minimization through a sliding window combinatorial optimization. By introducing principled metrics to quantify descriptor-space redundancy and information content, MSA selects a minimal set of keyframes that maintains loop closure potential and pose graph consistency, while remaining suitable for real-time application. (c) Through our results, we demonstrate that MSA provides higher localization accuracy in terms of absolute trajectory error and stronger place recognition performance across diverse environments, without manual parameter tuning. Compared to adaptive baselines such as entropy-based and spaciousness, MSA achieves state-of-the-art accuracy with comparable computational and memory demands, and compared to keeping all samples, it reduces loop closure query load and memory usage, making it well-suited for large-scale missions.
II Related Work
Efficient implementations of pose graph optimization have significantly enhanced its scalability by leveraging factor graphs and incremental optimization techniques. In more detail, GTSAM [gtsam] models probabilistic relationships using factor graphs and solves large-scale problems efficiently with solvers like iSAM2 [Kaess2011isam2], which incrementally updates the SLAM solution in real-time through a Bayes tree representation. Similarly, g2o [g2o] provides a versatile graph optimization framework, leveraging sparse pose graph connectivity for efficiency while supporting various parameterizations and loss functions. Furthermore, sparsification methods [bai2021sparsification, Nam2024spectral] and hierarchical approaches [Guadagnino2022hipe, Tazaki2022multires] have further improved efficiency by reducing the number of active nodes in the optimization process. Complementary to these back-end advances, Latif et al. [rrr2013] proposed RRR, which monitors pose-graph consistency over time and incrementally suppresses or removes spurious loop closures, improving robustness in long-term and multi-session mapping. However, despite these advances, loop closure detection remains a major bottleneck due to the large number of candidate pairs that must be checked and processed to add meaningful constraints to the pose graph.
Keyframe sampling helps reduce the computational demands of loop closure detection by limiting the number of frames stored and processed in SLAM pipelines. In LiDAR-(Inertial) Odometry, systems such as LIO-SAM [liosam2020shan] and DiSCo-SLAM [Huang2022disco] or back-end global optimization systems like LAMP [Chang2022lamp2], use fixed Euclidean distance intervals (1–2 meters) for keyframe generation to maintain local maps and optimize global poses. Although effective, these methods lack adaptability in dynamic environments with varying densities of loop closure candidates. Recent advancements have introduced adaptive keyframe sampling techniques, adjusting intervals based on environmental spaciousness [chen2022dlo, Kim2023adaptive], although manual threshold adjustment is still required. Entropy-based methods, like the one proposed by Zeng et al. [Zeng2023entropy], use information theory to select keyframes, but face similar adaptability issues. Other approaches, such as the one proposed by Ou et al. [Ou2021], focus on displacement vector similarity to balance computational cost and map completeness, while Chen et al. [degkey] select keyframes via online image-quality scoring (e.g., exposure, blur, occlusion), skipping degraded frames to improve visual localization. In summary, while prior works have made progress in improving pose graph optimization scalability and in proposing various keyframe sampling heuristics, they often treat loop closure detection, place recognition, and keyframe selection as separate problems. Our approach differs by jointly formulating these components within a unified framework, ensuring that sampling decisions directly account for their downstream impact on both loop closure detection and pose graph optimization.
III Problem Formulation
III-A Pose Graph Optimization
@@
The objective of PGO is to find the most likely configuration of a robot’s trajectory by minimizing the error in a graph-based representation of the poses. The pose graph is represented as a set of nodes and edges, where each node represents a robot pose and each edge encodes a spatial constraint between two poses [Grisetti2010graphslam]. Let $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ be the set of $N$ robot poses, where, $\mathbf{x}\in SE(3)$, $\mathbf{z}_{ij}=h(\mathbf{x}_{i},\mathbf{x}_{j})$ be the observed spatial constraint between a pair of poses $\langle i,j\rangle$ and $\hat{\mathbf{z}}_{ij}(\mathbf{x}_{i},\mathbf{x}_{j})$ be the predicted value of that constraint. The error $\mathbf{e}_{ij}\in\mathbb{R}^{6}$ quantifies the difference between the observed relative transformation and the predicted transformation between poses $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ and is denoted as:
$\mathbf{e}_{ij}=\mathbf{e}(\mathbf{x}_{i},\mathbf{x}_{j})=\mathbf{z}_{ij}-\hat{\mathbf{z}}_{ij}$ (1)
@@
We can then define the set of edges $\mathcal{E}=\big\{\langle i,j\rangle\>|\>\exists\,\mathbf{z}_{ij}\big\}$, that contains all the pairs in which a relative constraint exists. The goal of pose graph optimization is to find the optimal configuration of poses $\mathbf{X}^{*}$ that minimizes the overall error function $E(\mathbf{X})$, denoted as:
@@
III-B Odometry and Loop Closure Edges
Pose graph optimization is widely used as a global optimization method to mitigate accumulated pose drift during large-scale and long-term missions. This would not be possible without loop closures. Relying solely on pose-to-pose odometry constraints is insufficient to maintain global consistency, hence, loop closures are essential. Given this, we define two distinct subsets in the set of edges $\mathcal{E}$ in the pose graph. The first subset, $\mathcal{E}_{o}\subseteq\mathcal{E}$, consists of odometry edges, which represent consecutive movements of the robot based on local motion estimates, while the second subset, $\mathcal{E}_{l}\subseteq\mathcal{E}$, consists of loop closure edges, which correspond to observations of previously visited locations. Thus, the overall error function can be decomposed into two components:
@@ III-B Odometry and Loop Closure Edges
III-C Loop Closure Detection
Detecting loop closures is challenging, particularly in large-scale missions, where accumulated drift can cause traditional radius-based searches to fail by missing nearby candidates. A common solution is place recognition, which compares current sensory data (e.g., images or LiDAR scans) against a database of past observations to identify matches. To formalize this, we define a keyframe set $\mathbf{K}=\{(\mathbf{x}_{i},\mathbf{d}_{i},\mathbf{z}_{i})\,|\,i=1,\ldots,N\}$, where $\mathbf{x}_{i}\in SE(3)$ is the pose of keyframe $i$, $\mathbf{d}_{i}\in\mathbb{R}^{M}$ represents its descriptor which encodes distinct features [ma2022overlaptrans, Giseop2022scpp], and $\mathbf{z}_{i}$ is the set of associated measurements. The goal is to find loop closure candidates $\mathcal{L}$ by comparing descriptors:
$\mathcal{L}=\big\{\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}\times\mathbf{K},\;\mathbf{k}_{i}\neq\mathbf{k}_{j}\big\},$ (4)
@@ III-C Loop Closure Detection
III-D Reducing Search Space in Place Recognition
The detection and verification of loop closure edges $\mathcal{E}_{l}\triangleq\mathcal{L}$ is computationally demanding due to the large number of potential keyframe pairs, which scales quadratically as $O(n^{2})$, where $n=|\mathbf{K}|$ is the number of keyframes. While using descriptors and methods like k-nearest neighbor (k-NN) search [yin2023survey] can accelerate the process, additional steps are still required to remove outliers, mitigate false positives, or prioritize certain nodes [Denniston2022LoopCP], while extracting the relative pose transformation between loop closure nodes also involves computationally intensive techniques such as General Iterative Closest Point (GICP) [GICP]. This pipeline does not scale well for extensive missions, often resulting in processing times of 3-4 hours or more [Kamak2024darpa_slam, Chang2022lamp2], and the challenge becomes even greater in multi-robot systems [stathoulopoulos2024frame].
@@ III-D Reducing Search Space in Place Recognition
[…]
$\mathcal{L}^{*}=\big\{\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\mathbf{K}^{*}_{\text{S}}\times\mathbf{K}^{*}_{\text{S}}\big\},$ (6)
@@ III-E Underlying Challenges
Addressing the combined problem of pose graph optimization and loop closure detection through place recognition, presents several underlying challenges that affect both the efficiency and accuracy of the solution.
III-E1 Impact of Sampling
The sampling strategy $S$ reduces $\mathbf{K}$ to a smaller subset $\mathbf{K}_{\text{S}}$ to keep place recognition scalable, thereby altering the loop closure set and the odometry edges. Over-aggressive pruning can remove informative loop closures needed to correct drift, and over-coarsen the odometry chain, increasing linearization error and degrading PGO conditioning. It may also leave regions under-represented in descriptor space, weakening place recognition retrieval on revisit. Conversely, keeping many near-duplicate frames bloats place recognition search and verification, increases false positives in self-similar areas, and on the PGO side, over-weights correlated constraints, slowing or destabilizing the solver. The core challenge is to select a subset that preserves the essential loop closure structure while maintaining well-spaced consecutive poses, as depicted in Fig. 2.
III-E2 Combined Problem
Place recognition and pose graph optimization solve different subproblems. Place recognition operates in high-dimensional descriptor space, ranking appearance-based matches and remaining sensitive to viewpoint and self-similar structure; its errors show up as false positives or missed matches. PGO operates in geometric space on poses in $SE(3)$, solving a sparse nonlinear least-squares problem whose behavior depends on graph connectivity, scaling, and numerical conditioning; failures arise from inconsistent or poorly scaled constraints and weak observability. The challenge is to reconcile appearance-based data association with metric consistency over the pose graph.
III-E3 Dynamic Keyframe Selection